OpenStudy (xxorejina):

A company is manufacturing a gear in the shape of a regular polygon. The measure of each angle of the gear is 162º. How many sides does the gear have? A. 21 B. 18 C. 20 D. 19

OpenStudy (bibby):

the rule for the individual angle measurement of a (regular) n-sided polygon is \((n-2) × \dfrac{180}{n} \)

OpenStudy (bibby):

we're solving for n here, so we have \(162=(n-2)*\dfrac{180}{n}\\162n=180(n-2)\\162n=180n-360\\-18n=-360\\18n=360\\n=20\)

OpenStudy (xxorejina):

Thanks (: @bibby

OpenStudy (bibby):

np ;)